# Towards Adversarially Robust Dataset Distillation by Curvature Regularization

### v1

@@ Abstract
Dataset distillation (DD) allows datasets to be distilled to fractions of their original size while preserving the rich distributional information so that models trained on the distilled datasets can achieve a comparable accuracy while saving significant computational loads. Recent research in this area has been focusing on improving the accuracy of models trained on distilled datasets. In this paper, we aim to explore a new perspective of DD. We study how to embed adversarial robustness in distilled datasets, so that models trained on these datasets maintain the high accuracy and meanwhile acquire better adversarial robustness. We propose a new method that achieves this goal by incorporating curvature regularization into the distillation process with much less computational overhead than standard adversarial training. Extensive empirical experiments suggest that our method not only outperforms standard adversarial training on both accuracy and robustness with less computation overhead but is also capable of generating robust distilled datasets that can withstand various adversarial attacks.
Preprint. Under review.We will release our code upon acceptance.
1 Introduction
@@
In the era of big data, the computational demands for training deep learning models are continuously growing due to the increasing volume of data. This presents substantial challenges, particularly for entities with limited computational resources. To mitigate such issues, concepts like dataset distillation (Wang et al., 2018) and dataset condensation (Zhao et al., 2021; Zhao and Bilen, 2021, 2023) have emerged, offering a means to reduce the size of the data while maintaining its utility. A successful implementation of dataset distillation can bring many benefits, such as enabling more cost-effective research on large datasets and models.
@@
Dataset distillation refers to the task of synthesizing a smaller dataset such that models trained on this smaller set yield high performance when tested against the original, larger dataset. The dataset distillation algorithm takes a large dataset as input and generates a compact, synthetic dataset. The efficacy of the distilled dataset is assessed by evaluating models trained on it against the original dataset.
@@
Conventionally, distilled datasets are evaluated based on their standard test accuracy. Therefore, recent research has expanded rapidly in the direction of improving the test accuracy following the evaluation procedure (Sachdeva and McAuley, 2023). Additionally, many studies focus on improving the efficiency of the distillation process (Sachdeva and McAuley, 2023).
@@
Less attention, however, has been given to an equally important aspect of this area of research: the adversarial robustness of models trained on distilled datasets. Adversarial robustness is a key indicator of a model’s resilience against malicious inputs, making it a crucial aspect of trustworthy machine learning. Given the potential of dataset distillation to safeguard the privacy of the original dataset (Geng et al., 2023; Chen et al., 2023), exploring its capability to also enhance model robustness opens a promising avenue for advancing research in trustworthy machine learning (Liu et al., 2023a).
@@
Thus, our work seeks to bridge this gap and focuses on the following question: How can we embed adversarial robustness into the dataset distillation process, thereby generating datasets that lead to more robust models? Motivated by this question, we explore potential methods to accomplish this goal. As it turns out, it is not as simple as adding adversarial training to the distillation process. To find a more consistent method, we study the theoretical connection between adversarial robustness and dataset distillation. Our theory suggests that we can directly improve the robustness of the distilled dataset by minimizing the curvature of the loss function with respect to the real data. Based on our findings, we propose a novel method, GUARD (Geometric regUlarization for Adversarial Robust Dataset), which incorporates curvature regularization into the distillation process. We then evaluate GUARD against existing distillation methods on ImageNette, Tiny ImageNet, and ImageNet datasets. In summary, the contributions of this paper are as follows
@@
The remainder of this paper is structured as follows. In Section 2, we introduce a range of related works that provide context for our research. In Section 3, we offer a background to our study with a concise overview of dataset distillation, some preliminary results, and a formulation of the robust dataset distillation problem. In Section 4, we first explore the theoretical connections between dataset distillation and adversarial robustness, and then introduce the method, GUARD, based on our findings. The experiment results and discussions are subsequently detailed in Section 5 and Section 6, and we conclude the paper with a summary in Section 7.
2 Related Works
2.1 Dataset Distillation
@@
Dataset distillation is a technique that has been developed to address the issue of the increasing amount of data required to train deep learning models. The goal of dataset distillation is to efficiently train neural networks using a small set of synthesized training examples from a larger dataset. Dataset distillation (DD) (Wang et al., 2018) was one of the first such methods developed, and it showed that training on a few synthetic images can achieve similar performance on MNIST and CIFAR10 as training on the original dataset. Later, Cazenavette et al. (2022); Zhao and Bilen (2021); Zhao et al. (2021) explored different methods of distillation, including gradient and trajectory matching. These approaches focused on matching the gradient w.r.t. the real and synthetic data, with stronger supervision for the training process. Instead of matching the weights of the neural network, another thread of works (Lee et al., 2022; Wang et al., 2022) focuses on matching feature distributions of the real and synthetic data in the embedding space to better align features or preserve real-feature distribution. Considering the lack of efficiency of the bi-level optimization in previous methods, Nguyen et al. (2021); Zhou et al. (2022) aim to address the significant amount of meta gradient computation challenges. Nguyen et al. (2020) proposed a kernel-inducing points meta-learning algorithm and they further leverage the connection between the infinitely wide ConvNet and kernel ridge regression for better performance. Furthermore, Sucholutsky and Schonlau (2021) focuses on simultaneously distilling images and the paired soft labels. Last but not least, Yin et al. (2023) introduced SRe2L, which divided dataset distillation into three steps to make the process more efficient.
@@
In general, dataset distillation approaches can be broadly classified into four families based on their underlying principles: meta-model matching, gradient matching, distribution matching, and trajectory matching (Sachdeva and McAuley, 2023). Regardless of the particular approach, most of the existing methods rely on optimizing the distilled dataset w.r.t. a network trained with real data, such methods include DD (Wang et al., 2018), DC (Zhao et al., 2021), DSA (Zhao and Bilen, 2021), MTT (Cazenavette et al., 2022), DCC (Lee et al., 2022), SRe2(Yin et al., 2023), and many more.
2.2 Adversarial Attacks
@@
Adversarial attacks are a significant concern in the field of machine learning, as they can cause models to make incorrect predictions even when presented with seemingly similar input. Kurakin et al. (2017) demonstrates the real-world implications of these attacks. Many different types of adversarial attacks have been proposed in the literature (Goodfellow et al., 2015; Madry et al., 2018). In particular, Projected Gradient Descent (PGD) is a widely used adversarial attack that has been shown to be highly effective against a variety of machine learning models (Madry et al., 2018). The limitations of defensive distillation, a technique initially proposed for increasing the robustness of machine learning models, were explored by Papernot et al. . Moosavi-Dezfooli et al. (2016) introduced DeepFool, an efficient method to compute adversarial perturbations. Other notable works include the study of the transferability of adversarial attacks by Papernot et al. (2016a), the simple and effective black-box attack by Narodytska and Kasiviswanathan (2016), and the zeroth-order optimization-based attack by Chen et al. (2017). More recently, Athalye et al. (2018) investigated the robustness of obfuscated gradients, and Wong et al. (2019) introduced the Wasserstein smoothing as a novel defense against adversarial attacks. Croce and Hein (2020) introduced AutoAttack, which is a suite of adversarial attacks consisting of four diverse and parameter-free attacks that are designed to provide a comprehensive evaluation of a model’s robustness to adversarial attacks.
2.3 Adversarial Defense
@@
Numerous defenses against adversarial attacks have been proposed. Among these, adversarial training stands out as a widely adopted defense mechanism that entails training machine learning models on both clean and adversarial examples (Goodfellow et al., 2015). Several derivatives of the adversarial training approach have been proposed, such as ensemble adversarial training (Tramèr et al., 2018), and randomized smoothing (Cohen et al., 2019) — a method that incorporates random noise to obstruct the generation of effective adversarial examples. However, while adversarial training can be effective, it bears the drawback of being computationally expensive and time-consuming.
@@
Some defense mechanisms adopt a geometrical approach to robustness. One such defense mechanism is CURE, a method that seeks to improve model robustness by modifying the curvature loss function used during training (Moosavi-Dezfooli et al., 2019). The primary aim of CURE is to minimize the sensitivity of the model to adversarial perturbations in the input space to make it more difficult for an attacker to find adversarial examples which cross this boundary. Miyato et al. (2015) focuses on improving the smoothness of the output distribution to make models more resistant to adversarial attacks, while Cisse et al. (2017b) introduced Parseval networks, which enforce Lipschitz constant to improve model robustness. Ross and Doshi-Velez (2018) also presented a method for improving the robustness of deep learning models using input gradient regularization.
@@
Several other types of defense techniques have also been proposed, such as corrupting with additional noise and pre-processing with denoising autoencoders by Gu and Rigazio (2014), the defensive distillation approach by Papernot et al. (2016b), and the Houdini adversarial examples by Cisse et al. (2017a).
3 Preliminary
3.1 Dataset Distillation
@@
Let $\mathcal{T}$ denote the real dataset drawn from the distribution $\mathcal{D_{T}}$. $\mathcal{T}$ consists of $n$ image-label pairs $\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. Let $\mathcal{S}$ denote the distilled dataset drawn from the distribution $\mathcal{D_{S}}$, consisting of $m$ image-label pairs $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},\tilde{y}_{j})\}_{j=1}^{m}$. $\ell(\mathbf{x},y;\boldsymbol{\theta})$ denotes the loss function of a model parameterized by $\boldsymbol{\theta}$ on a sample $(\mathbf{x},y)$, and $\mathcal{L}(\mathcal{T};\boldsymbol{\theta})$ denotes the empirical loss on $\mathcal{T}$, $\mathcal{L}(\mathcal{T};\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{x}_{i},y_{i};\boldsymbol{\theta})$.
Given the real training set $\mathcal{T}$, Dataset Distillation (DD) aims at finding the optimal synthetic set $\mathcal{S}^{*}$ by solving a bi-level optimization problem as below:
@@
$\begin{gathered}\min_{\mathcal{S}}D({\boldsymbol{\theta}(\mathcal{S})},{\boldsymbol{\theta}(\mathcal{T})})\\ \textrm{subject to}\quad\boldsymbol{\theta}(\mathcal{S})=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathcal{S};\boldsymbol{\theta}),\;\;\boldsymbol{\theta}(\mathcal{T})=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathcal{T};\boldsymbol{\theta}),\end{gathered}$ (2)
where $D$ is a manually chosen distance function. Recent works have proliferated along this direction, with methods such as gradient matching (Zhao et al., 2021) and trajectory matching (Cazenavette et al., 2022), each focusing on aligning different aspects of the model’s optimization dynamics. The goal is to learn a distilled dataset that equips the model with properties similar to those imparted by the real dataset.
Some works have also tried to align the distribution of the distilled data with that of the real data (Zhao and Bilen, 2023; Zhang et al., 2024; Liu et al., 2023b), or recover a distilled version of the training data from a trained model (Yin et al., 2023; Buzaglo et al., 2023). These methods do not rely on the computation of second-order gradients, leading to improved efficiency and performance on large-scale datasets.
@@
3.2 The Limitation of Adversarial Training in Dataset Distillation
@@
In the supervised learning setting, one of the most commonly used methods to enhance model robustness is adversarial training, which involves training the model on adversarial examples that are algorithmically searched for or crafted (Goodfellow et al., 2015). This can be formulated as
@@
Analogously, in the dataset distillation setting, one intuitive way to distill robust datasets would be to synthesize a distilled dataset using a robust model trained with adversarial training. As mentioned in Section 2, many dataset distillation methods utilize a model trained on the original dataset as a comparison target, therefore this technique can be easily integrated to those methods.
While embedding adversarial training directly within the dataset distillation process may seem like an intuitive and straightforward approach, our comprehensive analysis reveals its limitations across various distillation methods. As an example, we show the evaluation of one such implementation using SRe2L, a recently proposed efficient dataset distillation method in Table 1. The results indicate a significant decline in clean accuracy for models trained on datasets distilled using this technique, in contrast to those synthesized by the original method. Moreover, the improvements in robustness achieved are very inconsistent. In our experiment, we only employed a weak PGD attack with $\epsilon=1/255$ to generate adversarial examples for adversarial training, leading to the conclusion that even minimal adversarial training can detrimentally impact model performance when integrated into the dataset distillation process.
@@
4 Methods
4.1 Formulation of the Robust Dataset Distillation Problem
Following our notations in Section 3.1, and extending the distillation problem to the adversarial robustness setting, robust dataset distillation can be formulated as a tri-level optimization problem as below:

### v4

@@ Abstract
Dataset distillation (DD) allows datasets to be distilled to fractions of their original size while preserving the rich distributional information, so that models trained on the distilled datasets can achieve a comparable accuracy while saving significant computational loads. Recent research in this area has been focusing on improving the accuracy of models trained on distilled datasets. In this paper, we aim to explore a new perspective of DD. We study how to embed adversarial robustness in distilled datasets, so that models trained on these datasets maintain the high accuracy and meanwhile acquire better adversarial robustness. We propose a new method that achieves this goal by incorporating curvature regularization into the distillation process with much less computational overhead than standard adversarial training. Extensive empirical experiments suggest that our method not only outperforms standard adversarial training on both accuracy and robustness with less computation overhead but is also capable of generating robust distilled datasets that can withstand various adversarial attacks. Our implementation is available at: https://github.com/yumozi/GUARD.
Introduction
@@
In the era of big data, the computational demands for training deep learning models are continuously growing due to the increasing volume of data. This presents substantial challenges, particularly for entities with limited computational resources. To mitigate such issues, concepts like dataset distillation (Wang et al. 2018) and dataset condensation (Zhao, Mopuri, and Bilen 2021; Zhao and Bilen 2021, 2023) have emerged, offering a means to reduce the size of the data while maintaining its utility. A successful implementation of dataset distillation can bring many benefits, such as enabling more cost-effective research on large datasets and models.
@@
Dataset distillation (DD) refers to the task of synthesizing a smaller dataset such that models trained on this smaller set yield high performance when tested against the original, larger dataset. Dataset distillation algorithms take a large dataset as input and generate a compact, synthetic dataset. The efficacy of the distilled dataset is assessed by evaluating models trained on it against the original dataset.
@@
Conventionally, distilled datasets are evaluated based on their standard test accuracy. Therefore, recent research has expanded rapidly in the direction of improving the test accuracy following the evaluation procedure (Sachdeva and McAuley 2023). Additionally, many studies focus on improving the efficiency of the distillation process (Sachdeva and McAuley 2023).
@@
Less attention, however, has been given to an equally important aspect of this area of research: the adversarial robustness of models trained on distilled datasets. Adversarial robustness is a key indicator of a model’s resilience against malicious inputs, making it a crucial aspect of trustworthy machine learning. Given the potential of dataset distillation to safeguard the privacy of the original dataset (Geng et al. 2023; Chen et al. 2023), exploring its capability to also enhance model robustness opens a promising avenue for advancing research in trustworthy machine learning (Liu, Chaudhary, and Wang 2023). Thus, our work seeks to bridge this gap and focuses on the following question: How can we embed adversarial robustness into the dataset distillation process, thereby generating datasets that lead to more robust models?
@@
Motivated by this question, we explore potential methods to accomplish this goal. As it turns out, it is not as simple as adding adversarial training to the distillation process. To find a more consistent method, we study the theoretical connection between adversarial robustness and dataset distillation. Our theory suggests that we can directly improve the robustness of the distilled dataset by minimizing the curvature of the loss function with respect to the real data. Based on our findings, we propose a novel method, GUARD (Geometric Regularization for Adversarially Robust Dataset), which incorporates curvature regularization into the distillation process. We then evaluate GUARD against existing distillation methods on ImageNette, Tiny ImageNet, and ImageNet datasets. In summary, the contributions of this paper are as follows
@@
Related Works
Dataset Distillation
@@
Aiming to address the issue of the increasing amount of data required to train deep learning models, the goal of dataset distillation is to efficiently train neural networks using a small set of synthesized training examples from a larger dataset. Dataset distillation (DD) (Wang et al. 2018) was one of the first such methods developed, and it showed that training on a few synthetic images can achieve similar performance on MNIST and CIFAR10 as training on the original dataset. Later, Cazenavette et al. (2022); Zhao and Bilen (2021); Zhao, Mopuri, and Bilen (2021); Lee et al. (2022) explored different methods of distillation, including gradient and trajectory matching w.r.t. the real and synthetic data, with stronger supervision for the training process. Instead of matching the weights of the neural network, another thread of works (Wang et al. 2022; Zhao and Bilen 2023; Zhang et al. 2024; Liu et al. 2023) focuses on matching feature distributions of the real and synthetic data in the embedding space to better align features or preserve real-feature distribution. Considering the lack of efficiency of the bi-level optimization in previous methods, Nguyen et al. (2021); Zhou, Nezhadarya, and Ba (2022) aim to address the significant amount of meta gradient computation challenges. Nguyen, Chen, and Lee (2020) proposed a kernel-inducing points meta-learning algorithm and they further leverage the connection between the infinitely wide ConvNet and kernel ridge regression for better performance. Furthermore, Sucholutsky and Schonlau (2021) addresses the simultaneous distillation of images and their corresponding soft labels. Later, some works focused on further improving efficiency of the process, such as Yin, Xing, and Shen (2023) that introduced SRe2L, which optimizes the distillation process by dividing it into three distinct steps for greater efficiency, and Xu et al. (2024), which proposed an approach to enhance both the efficiency and performance by first pruning the original dataset. Finally, Li et al. (2024) further advanced the process by dynamically pruning the original dataset based on the desired compression ratio and extracting information from deeper layers of the network.
@@
Dataset distillation approaches can be broadly classified into four families based on their underlying principles: meta-model matching, gradient matching, distribution matching, and trajectory matching (Sachdeva and McAuley 2023). Regardless of the particular approach, most of the existing methods rely on optimizing the distilled dataset w.r.t. a network trained with real data, such methods include DD (Wang et al. 2018), DC (Zhao, Mopuri, and Bilen 2021), DSA (Zhao and Bilen 2021), MTT (Cazenavette et al. 2022), DCC (Lee et al. 2022), SRe2(Yin, Xing, and Shen 2023), ATT (Liu et al. 2024) and many more.
In a related direction, some works also address the robustness of dataset distillation, but specifically focusing on out-of-distribution (OOD) robustness. For instance, Vahidian et al. (2024) employs risk minimization techniques to ensure robustness, while TrustDD (Ma et al. 2024) incorporates outliers during the distillation process to facilitate OOD detection.
Adversarial Attacks
@@
Adversarial attacks are a significant concern in the field of machine learning, as they can cause models to make incorrect predictions even when presented with seemingly similar input. Kurakin, Goodfellow, and Bengio (2017) demonstrates the real-world implications of these attacks. Many different types of adversarial attacks have been proposed in the literature (Goodfellow, Shlens, and Szegedy 2015; Madry et al. 2018). In particular, Projected Gradient Descent (PGD) is a widely used adversarial attack that has been shown to be highly effective against a variety of machine learning models (Madry et al. 2018). The limitations of defensive distillation, a technique initially proposed for increasing the robustness of machine learning models, were explored by Papernot et al. (2017). Moosavi-Dezfooli, Fawzi, and Frossard (2016) introduced DeepFool, an efficient method to compute adversarial perturbations. Other notable works include the study of the transferability of adversarial attacks by Papernot, McDaniel, and Goodfellow (2016), the simple and effective black-box attack by Narodytska and Kasiviswanathan (2016), and the zeroth-order optimization-based attack by Chen et al. (2017). More recently, Athalye, Carlini, and Wagner (2018) investigated the robustness of obfuscated gradients, and Wong, Schmidt, and Kolter (2019) introduced the Wasserstein smoothing as a novel defense against adversarial attacks. Croce and Hein (2020) introduced AutoAttack, which is a suite of adversarial attacks consisting of four diverse and parameter-free attacks that are designed to provide a comprehensive evaluation of a model’s robustness to adversarial attacks.
Adversarial Defense
@@
Numerous defenses against adversarial attacks have been proposed. Among these, adversarial training stands out as a widely adopted defense mechanism that entails training machine learning models on both clean and adversarial examples (Goodfellow, Shlens, and Szegedy 2015). Several derivatives of the adversarial training approach have been proposed, such as ensemble adversarial training (Tramèr et al. 2018), and randomized smoothing (Cohen, Rosenfield, and Kolter 2019). However, while adversarial training can be effective, it bears the drawback of being computationally expensive and time-consuming.
@@
Some defense mechanisms adopt a geometrical approach to robustness. One such defense mechanism is CURE  (Moosavi-Dezfooli et al. 2019), a method that seeks to improve model robustness by reducing the curvature of the loss landscape during training. Similarly, Miyato et al. (2015) improved the smoothness of the output distribution, Cisse et al. (2017b) enforced Lipschitz constants, Ross and Doshi-Velez (2018) employed input gradient regularization, to improve the models’ adversarial robustness.
@@
Several other types of defense techniques have also been proposed, such as corrupting with additional noise and pre-processing with denoising autoencoders by Gu and Rigazio (2014), the defensive distillation approach by Papernot et al. (2016), the Houdini adversarial examples by Cisse et al. (2017a), and the approximate null space augmentation by Liu et al. (2025).
Preliminary
Dataset Distillation
@@
Let $\mathcal{T}$ represent the real dataset, drawn from the distribution $\mathcal{D_{T}}$. The dataset $\mathcal{T}$ comprises $n$ image-label pairs, defined as $\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. Similarly, let $\mathcal{S}$ denote the distilled dataset, drawn from the distribution $\mathcal{D_{S}}$, and consisting of $m$ image-label pairs, defined as $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},\tilde{y}_{j})\}_{j=1}^{m}$, where $m\ll n$. Conventionally, instead of directly expressing the size of the distilled dataset as $|\mathcal{S}|$, it is more common to describe it in terms of “images per class” (IPC). Let $\ell(\mathbf{x},y;\boldsymbol{\theta})$ denote the loss function of a model parameterized by $\boldsymbol{\theta}$ on a sample $(\mathbf{x},y)$, and $\mathcal{L}(\mathcal{T};\boldsymbol{\theta})$ denotes the empirical loss on $\mathcal{T}$, $\mathcal{L}(\mathcal{T};\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{x}_{i},y_{i};\boldsymbol{\theta})$.
Given the real training set $\mathcal{T}$, dataset distillation aims to find the optimal synthetic dataset $\mathcal{S}^{*}$ by solving the following bi-level optimization problem:
@@
$\begin{gathered}\min_{\mathcal{S}}D({\boldsymbol{\theta}(\mathcal{S})},{\boldsymbol{\theta}(\mathcal{T})})\\ \textrm{subject to}\quad\boldsymbol{\theta}(\mathcal{S})=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathcal{S};\boldsymbol{\theta})\\ \textrm{and}\quad\boldsymbol{\theta}(\mathcal{T})=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathcal{T};\boldsymbol{\theta}),\end{gathered}$ (2)
where $D$ is a manually chosen distance function. Recent works have proliferated along this direction, with methods such as gradient matching (Zhao, Mopuri, and Bilen 2021) and trajectory matching (Cazenavette et al. 2022), each focusing on aligning different aspects of the model’s optimization dynamics. Some works have also tried to align the distribution of the distilled data with that of the real data (Zhao and Bilen 2023; Zhang et al. 2024; Liu et al. 2023), or recover a distilled version of the training data from a trained model (Yin, Xing, and Shen 2023; Buzaglo et al. 2023). These methods do not rely on the computation of second-order gradients, leading to improved efficiency and performance on large-scale datasets.
@@
The Limitation of Adversarial Training in Dataset Distillation
@@
In the supervised learning setting, one of the most commonly used methods to enhance model robustness is adversarial training, which involves training the model on adversarial examples that are algorithmically searched for or crafted (Goodfellow, Shlens, and Szegedy 2015). This can be formulated as
@@
Analogously, in the dataset distillation setting, one intuitive way to distill robust datasets would be to synthesize a distilled dataset using a robust model trained with adversarial training. As mentioned in the related works section, many dataset distillation methods utilize a model trained on the original dataset as a comparison target, therefore this technique can be easily integrated to those methods.
While embedding adversarial training directly within the dataset distillation process may seem like an intuitive and straightforward approach, our comprehensive analysis reveals its limitations across various distillation methods. As an example, we show the evaluation of one such implementation based on SRe2(Yin, Xing, and Shen 2023) in Table 1. The results indicate a significant decline in clean accuracy for models trained on datasets distilled using this technique, in contrast to those synthesized by the original method. Moreover, the improvements in robustness achieved are very inconsistent. In our experiment, we only employed a weak PGD attack with $\epsilon=1/255$ to generate adversarial examples for adversarial training, leading to the conclusion that even minimal adversarial training can detrimentally impact model performance when integrated into the dataset distillation process.
@@
Methods
Formulation of the Robust Distillation Problem
Extending the distillation problem to the adversarial robustness setting, robust dataset distillation can be formulated as a tri-level optimization problem as below: